# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ 2.2. Intersection of block subspaces
Since $z_{n_{i}}\in\langle A\rangle\cap\langle B\rangle$ for all $i$, we have that $\langle C\rangle\subseteq\langle A\rangle\cap\langle B\rangle$. Conversely, let $v\in\langle A\rangle\cap\langle B\rangle$. Let $n_{i_{0}},\dots,n_{i_{t-1}}$ enumerate all $n_{i_{j}}$’s such that $n_{i_{j}}\in\operatorname{supp}_{A}(v)$. Fix any $k_{j}\in\operatorname{supp}(x_{n_{i_{j}}})\subseteq\operatorname{supp}(z_{n_{i_{j}}})$, and let:
$\lambda_{j}:=\frac{v(k_{j})}{z_{n_{i_{j}}}(k_{j})}.$
Note that $\lambda_{j}\neq 0$ for all $j$. Now let:
$v^{\prime}:=\sum_{j<t}\lambda_{j}z_{n_{i_{j}}}.$
@@
We first note that $\operatorname{supp}(v^{\prime})=\operatorname{supp}(v)$ – observe that $\operatorname{supp}_{A}(v^{\prime})=\bigcup_{j<t}\operatorname{supp}(z_{n_{i_{j}}}$). If $m\in\operatorname{supp}_{A}(v)$, then $m]_{\sim}\subseteq\operatorname{supp}_{A}(v)$ and $m]_{\sim}=[n_{i_{j}}]_{\sim}=\operatorname{supp}_{A}(z_{n_{i_{j}}})$ for some $j$, so $m\in\operatorname{supp}_{A}(v^{\prime})$. Conversely, if $m\in\operatorname{supp}(z_{n_{i_{j}}})$, then since $n_{i_{j}}\in\operatorname{supp}_{A}(v)$, we have that $m]_{\sim}=[n_{i_{j}}]_{\sim}\subseteq\operatorname{supp}_{A}(v)$, so $m\in\operatorname{supp}(v)$.
It remains to show that $v^{\prime}(k)=v(k)$ for all $k\in\operatorname{supp}(v)$. We fix some $j<t$, and observe that:
$v^{\prime}(k_{j})=\frac{v(k_{j})}{z_{n_{i_{j}}}(k_{j})}\cdot z_{n_{i_{j}}}(k_{j})=v(k_{j}),$
Since $v\in\langle A\rangle$ and $k_{j}\in\operatorname{supp}(x_{n_{i_{j}}})\subseteq\operatorname{supp}(v)$, there exist some vectors $u,w\in\langle A\rangle$ such that $u<x_{n_{i_{j}}}<w$, and:
$v=u+\frac{v(k_{j})}{x_{n_{i_{j}}}(k_{j})}x_{n_{i_{j}}}+w.$
Similarly, there exist some vectors $u^{\prime},w^{\prime}\in\langle A\rangle$ such that $u^{\prime}<x_{n_{i_{j}}}<w^{\prime}$, and:
$v^{\prime}=u^{\prime}+\frac{v(k_{j})}{x_{n_{i_{j}}}(k_{j})}x_{n_{i_{j}}}+w^{\prime}.$
Therefore, $v-v^{\prime}\in\langle A\rangle\cap\langle B\rangle$, and $n_{i_{j}}\notin\operatorname{supp}(v-v^{\prime})$. But this implies that $n_{i_{j}}]_{\sim}\cap\operatorname{supp}(v-v^{\prime})=\emptyset$, which is only possible if for all $k\in\operatorname{supp}(z_{n_{i_{j}}})$, $v^{\prime}(k)=v(k)$. Therefore, for all $j<t$ and $k\in\operatorname{supp}(z_{n_{i_{j}}})$, $v^{\prime}(k)=v(k)$, so $v^{\prime}=v$, as desired. ∎
@@ 3. Local Ramsey theory of block subspaces
Smythe explored the local Ramsey theory of block subspaces of a countable vector space developed in [S18], mainly focusing on $\leq^{*}$-upward closed subsets of $E^{[\infty]}$ (which Smythe calls families) and their various properties. We provide an overview of this theory, along with several additional results which will prove to be useful in later sections.
@@ 3.1. Semicoideals
To cope with the failure of the pigeonhole principle in the setting of countable vector spaces with $|\mathbb{F}|\geq 3$, Smythe introduced a weaker form of the pigeonhole property of semicoideals, which he calls fullness.
@@
It turns out that if $|\mathbb{F}|=2$, then fullness coincides with the local pigeonhole principle property. In this case, we say that $\mathcal{H}$ is a coideal.
@@
By Hindman’s theorem, $E^{[\infty]}$ is a coideal if $|\mathbb{F}|=2$.
@@
Suppose that $|\mathbb{F}|=2$, and let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. Then $\mathcal{H}$ is full iff $\mathcal{H}$ is a coideal.
@@
Suppose that $\mathcal{H}$ is a coideal. Let $Y\subseteq E$, and suppose that $Y$ is $\mathcal{H}$-dense below some $A\in\mathcal{H}$. Since $\mathcal{H}$ is a coideal, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$. Since $\mathcal{Y}$ is $\mathcal{H}$-dense below $A$, the latter option is not possible.
@@ 3.2. Selectivity properties
In the process of studying mad families (particularly $\mathcal{H}(\mathcal{A})$, see Definition 4.7), Smythe extended several local Ramsey theoretic properties of coideals of $\mathbb{N}$$^{\infty}$ to which Smythe extended to semicoideals. The first such property extends the (p)-property for coideals of $\mathbb{N}$$^{\infty}$ (see Lemma 7.4, [T10]).
@@
    Let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$. We say that $B\in\mathcal{H}$ weakly diagonalises $(A_{n})_{n<\omega}$ if $B\leq A_{0}$ and $B\leq^{*}A_{n}$ for all $n$.
@@
We shall provide a similar extension of the (q)-property (see Lemma 7.4, [T10]). Smythe also made a similar extension, which he termed as spreadness (Definition 8.11, [S18]).
@@
In a similar vein, we extend the definition of selectivity of coideals of $\mathbb{N}$$^{\infty}$ (see Definition 7.3, [T10]) to semicoideals of block subspaces.
Definition 3.8.
For any $A\in E^{[\infty]}$ and $x\in E$, we write:
$d_{A}(x):=\min\{N<\omega:\max(\operatorname{supp}(x))<A/N\}.$
@@
    Let $A\in\mathcal{H}$, and let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$ below $A$ (i.e. $A_{0}\leq A$). We say that $B\in\mathcal{H}\mathord{\upharpoonright}A$ diagonalises $(A_{n})_{n<\omega}$ within $A$ if $B/n\leq A_{d_{A}(r_{n}(B))}$ for all $n<\omega$.
@@
    We say that $\mathcal{H}$ is selective if for all $A\in\mathcal{H}$ and $\leq$-decreasing chain $(A_{n})_{n<\omega}$ of elements in $\mathcal{H}$ below $A$, there exists a diagonalisation of $(A_{n})_{n<\omega}$ within $A$ in $\mathcal{H}$.
@@
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a (q)-semicoideal. Let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$. If $B\in\mathcal{H}\mathord{\upharpoonright}A_{0}$ weakly diagonalises $(A_{n})_{n<\omega}$, then there exists some $C\in\mathcal{H}\mathord{\upharpoonright}B$ which diagonalises $(A_{n})_{n<\omega}$ within $A_{0}$.
@@
Let $A:=A_{0}$, and let $B\in\mathcal{H}\mathord{\upharpoonright}A$ weakly diagonalise $(A_{n})_{n<\omega}$. Let $(I_{m})_{m<\omega}$ be any increasing sequence of intervals such that $B/\max(I_{m})\leq A_{d_{A}(r_{\min(I_{m})-1}(B))}$. By the (q)-property, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}B$ such that for all $n$, $\langle r_{n}(C)\rangle\subseteq\langle r_{\min(I_{m})-1}(B)\rangle$ and $C/n\leq B/\max(I_{m})$. The first property implies that $d_{A}(r_{n}(C))\leq d_{A}(r_{\min(I_{m})}(B))$, so by the second property:
@@
so $C$ strongly diagonalises $(A_{n})_{n<\omega}$. ∎
@@
$\langle r_{n}(B)\rangle\subseteq\langle r_{k}(A)\rangle\subseteq\langle r_{\min(I_{m(k)})}(A)\rangle,$
@@
4. Almost disjoint families
@@ 4.1. Uncountability of almost disjoint families
We begin with the most fundamental question of the theory of mad families – the uncountability of infinite mad families. Consider the following two cardinals:
@@
It is clear that $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}\leq\mathfrak{c}$. By Proposition 2.5 of [S19], $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ (in fact, $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$), so $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ is a cardinal invariant. One may ask if the same holds for $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ - that is, if $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ is true. Smythe’s proof of $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ heavily relies on the following lemma:
Lemma 4.3 (Lemma 2.3, [S18]).
Let $A\in E^{[\infty]}$ be a block subspace, and let $a\in E^{[<\infty]}$ be a finite-dimensional block subspace. Then there exists some $N<\omega$, that depends only on $A$ and $\max(\operatorname{supp}(a))$, such that for all $x>N$ such that $x\notin\langle A\rangle$,
@@
He remarked that removing the “block” requirement from this lemma seems difficult. $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ was also stated as Proposition 3.1 of [K00], but Kolman further assumed that in the countable almost disjoint family of subspaces $\{V_{n}:n<\omega\}$, $V_{n}\cap\sum_{i\in I}V_{i}$ is finite-dimensional for all finite $n\notin I\subseteq\omega$. This is generally not satisfied by families of subspaces that are pairwise almost disjoint. The author raised this question on MathOverflow, of which user527492 provided an affirmative answer (see [u25]).
@@
Let $U,V,W\subseteq E$ be subspaces such that $\dim(V\cap W)<\infty$ and $\dim(U)<\infty$. Then $\dim((U+V)\cap W)<\infty$.
@@
Define a linear map $\phi:V\times W\to E$ by stipulating that $\phi(v,w):=w-v$. Let $\rho:E\to E/U$ be the canonical quotient map, and let $\psi:=\rho\circ\pi$. Note that:
@@
Let $\pi:V\oplus W\to W$ be the projection map. Observe that for all $w\in W$, we have that:
@@
We shall define the subspace $W=\{x_{n}:n<\omega\}$ inductively as follows: Let $x_{0}\in V_{0}$ be any vector. Now assume that we have defined a linearly independent set $\{x_{i}:i<n\}$ such that for all $k<n$, we have that:
$\operatorname{span}\{x_{i}:i<n\}\cap V_{k}\subseteq\operatorname{span}\{x_{i}:i\leq k\}.$
We wish to find some vector $x_{n}\in V_{n}$ such that for all $k<n$, we have that:
$\operatorname{span}\{x_{i}:i\leq n\}\cap V_{k}\subseteq\operatorname{span}\{x_{i}:i\leq k\}.$
If we can complete this induction step, then we are done as we have that $W\cap V_{n}\subseteq\operatorname{span}\{x_{0},\dots,x_{n}\}$ is finite-dimensional for all $n$.
For each $k<n$, let $U_{k}:=V_{k}+\operatorname{span}\{x_{i}:i<n\}$. By Claim 5, $V_{n}\cap U_{k}$ is of finite dimension. Since every infinite-dimensional vector space is not a finite union of finite-dimensional subspaces, we have that:
@@
so we may let $x_{n}$ be any vector in this set. We shall show that this vector $x_{n}$ works. We first note that $\{x_{i}:i\leq n\}$ is linearly independent, as $\operatorname{span}\{x_{i}:i<n\}\subseteq U_{0}$ and $x_{n}\notin U_{0}$. Now let $k<n$ and $x\in\operatorname{span}\{x_{i}:i<n+1\}\cap V_{k}$, so we have that $x=\sum_{i\leq n}\lambda_{i}x_{i}\in V_{k}$. If $\lambda_{n}\neq 0$, then we have that:
$x_{n}=\lambda_{n}^{-1}\left(x-\sum_{i<n}\lambda_{i}x_{i}\right)\in V_{n}\cap U_{k},$
contradicting the choice of $x_{n}$. Therefore, $\lambda_{n}=0$, which implies that:
$x\in\operatorname{span}\{x_{i}:i<n\}\cap V_{k}=\operatorname{span}\{x_{i}:i\leq k\}\cap V_{k},$
@@
Smythe asked in [S19] whether $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}=\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ holds in $\mathsf{ZFC}$, and this problem remains open to the author’s knowledge. In the same paper, Smythe showed that every finite almost disjoint family of block subspaces of cardinality greater than $1$ is not maximal. We extend this fact to arbitrary subspaces.
@@
Let $\mathcal{A}$ be a finite almost disjoint family of infinite-dimensional subspaces of $E$ such that $|\mathcal{A}|>1$. Then $\mathcal{A}$ is not maximal.
@@
Let $\mathcal{A}=\{V_{k}:k<n\}$ be a finite almost disjoint family of infinite-dimensional subspaces of $E$.
@@
We first assume that for all $j\neq k$, $V_{j}\cap V_{k}=\{0\}$. If $\mathbb{F}$ is infinite, then we are done, as any subspace over an infinte field is not a finite union of proper subspaces. Assume that $|\mathbb{F}|=N<\infty$, and suppose for a contradiction that $E=\bigcup_{k<n}V_{k}$. Let $M$ be some large number (which we shall fix later), and let $W:=\operatorname{span}\{e_{n}:n<M\}$. Let $U_{k}:=W\cap V_{k}$. Note that $W=\bigcup_{j<n}U_{k}$.

### v5

@@ 2.2. Intersection of block subspaces
Since $z_{n_{i}}\in\langle A\rangle\cap\langle B\rangle$ for all $i$, we have that $\langle C\rangle\subseteq\langle A\rangle\cap\langle B\rangle$. Conversely, let $v\in\langle A\rangle\cap\langle B\rangle$. Let $n_{i_{0}},\dots,n_{i_{t-1}}$ be an increasing enumeration of all $n_{i_{j}}$’s such that $n_{i_{j}}\in\operatorname{supp}_{A}(v)$. Fix any $k_{j}\in\operatorname{supp}(x_{n_{i_{j}}})\subseteq\operatorname{supp}(z_{n_{i_{j}}})$, and let:
$\xi_{j}:=\frac{\lambda_{k_{j}}(v)}{\lambda_{k_{j}}(z_{n_{i_{j}}})}.$
Note that $\xi_{j}\neq 0$ for all $j$. Now let:
$v^{\prime}:=\sum_{j<t}\xi_{j}z_{n_{i_{j}}}.$
@@
We first note that $\operatorname{supp}_{A}(v^{\prime})=\operatorname{supp}_{A}(v)$ (which implies that $\operatorname{supp}(v^{\prime})=\operatorname{supp}(v)$) — observe that $\operatorname{supp}_{A}(v^{\prime})=\bigcup_{j<t}\operatorname{supp}_{A}(z_{n_{i_{j}}})$. If $m\in\operatorname{supp}_{A}(v)$, then $m]_{\sim}\subseteq\operatorname{supp}_{A}(v)$ and $m]_{\sim}=[n_{i_{j}}]_{\sim}=\operatorname{supp}_{A}(z_{n_{i_{j}}})$ for some $j$, so $m\in\operatorname{supp}_{A}(v^{\prime})$. Conversely, if $m\in\operatorname{supp}_{A}(z_{n_{i_{j}}})$, then since $n_{i_{j}}\in\operatorname{supp}_{A}(v)$, we have that $m]_{\sim}=[n_{i_{j}}]_{\sim}\subseteq\operatorname{supp}_{A}(v)$, so $m\in\operatorname{supp}_{A}(v)$.
It remains to show that $\lambda_{k}(v^{\prime})=\lambda_{k}(v)$ for all $k\in\operatorname{supp}(v)$. We fix some $j<t$, and observe that:
$\lambda_{k_{j}}(v^{\prime})=\frac{\lambda_{k_{j}}(v)}{\lambda_{k_{j}}(z_{n_{i_{j}}})}\cdot\lambda_{k_{j}}(z_{n_{i_{j}}})=\lambda_{k_{j}}(v).$
Since $v\in\langle A\rangle$ and $k_{j}\in\operatorname{supp}(x_{n_{i_{j}}})\subseteq\operatorname{supp}(v)$, there exist some vectors $u,w\in\langle A\rangle$ (or $u=0$ or $w=0$) such that $u<x_{n_{i_{j}}}<w$, and:
$v=u+\frac{\lambda_{k_{j}}(v)}{\lambda_{k_{j}}(x_{n_{i_{j}}})}x_{n_{i_{j}}}+w.$
Similarly, there exist some vectors $u^{\prime},w^{\prime}\in\langle A\rangle$ such that $u^{\prime}<x_{n_{i_{j}}}<w^{\prime}$ (or $u^{\prime}=0$ or $w^{\prime}=0$), and:
$v^{\prime}=u^{\prime}+\frac{\lambda_{k_{j}}(v)}{\lambda_{k_{j}}(x_{n_{i_{j}}})}x_{n_{i_{j}}}+w^{\prime}.$
Therefore, $v-v^{\prime}\in\langle A\rangle\cap\langle B\rangle$, and $n_{i_{j}}\notin\operatorname{supp}(v-v^{\prime})$. But this implies that $n_{i_{j}}]_{\sim}\cap\operatorname{supp}(v-v^{\prime})=\emptyset$, which is only possible if for all $k\in\operatorname{supp}(z_{n_{i_{j}}})$, $\lambda_{k}(v^{\prime})=\lambda_{k}(v)$. Therefore, for all $j<t$ and $k\in\operatorname{supp}(z_{n_{i_{j}}})$, $\lambda_{k}(v^{\prime})=\lambda_{k}(v)$, so $v^{\prime}=v$, as desired. ∎
@@ 3. Local Ramsey theory of block subspaces
Smythe explored the local Ramsey theory of block subspaces of a countable vector space in [23], mainly focusing on $\leq^{*}$-upward closed subsets of $E^{[\infty]}$ (which Smythe calls families) and their various properties. We provide an overview of this theory, along with several additional results which will prove to be useful in later sections.
@@ 3.1. Semicoideals
To cope with the failure of the pigeonhole principle in the setting of countable vector spaces over $\mathbb{F}\neq\mathbb{F}_{2}$, Smythe introduced a weaker form of the pigeonhole property of semicoideals, which he calls fullness.
@@
It turns out that if $E$ is a vector space over $\mathbb{F}_{2}$, then fullness coincides with the local pigeonhole principle property. In this case, we say that $\mathcal{H}$ is a coideal.
@@
By Hindman’s theorem, $E^{[\infty]}$ is a coideal if $E$ is a vector space over $\mathbb{F}_{2}$.
@@
Suppose that $E$ is a vector space over $\mathbb{F}_{2}$, and let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. Then $\mathcal{H}$ is full iff $\mathcal{H}$ is a coideal.
@@
Suppose that $\mathcal{H}$ is a coideal. Let $Y\subseteq E$, and suppose that $Y$ is $\mathcal{H}$-dense below some $A\in\mathcal{H}$. Since $\mathcal{H}$ is a coideal, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$. Since $Y$ is $\mathcal{H}$-dense below $A$, the latter option is not possible.
@@ 3.2. Selectivity properties
In the process of studying semicoideals related to mad families (particularly $\mathcal{H}(\mathcal{A})$, see Definition 4.9), Smythe extended several local Ramsey theoretic properties of coideals of $\mathbb{N}$$^{\infty}$ to semicoideals of $E^{[\infty]}$. Selectivity properties of (ultra)filters (see Definition 7.5) in $E^{[\infty]}$ has also been studied in [25]. The first such property extends the (p)-property for coideals of $\mathbb{N}$$^{\infty}$ (see Lemma 7.4, [27]).
@@
    Let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$. We say that $B\in\mathcal{H}$ weakly diagonalises $(A_{n})_{n<\omega}$ if $B\leq A_{0}$ and $B\leq^{*}A_{n}$ for all nn [^3] In Definition 2.2 of [23], this property was similarly defined using the term diagonalises. We shall reserve this term for a stronger form of diagonalisation later (Definition 3.8).
@@
We shall provide a similar extension of the (q)-property (see Lemma 7.4, [27]). Smythe also made a similar extension, which he termed as spreadness (Definition 8.11, [23]).
@@
In a similar vein, we extend the definition of selectivity of coideals of $\mathbb{N}$$^{\infty}$ (see Definition 7.3, [27]) to semicoideals of block subspaces.
@@
    Let $A\in\mathcal{H}$, and let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$ below $A$ (i.e. $A_{0}\leq A$). We say that $B\in\mathcal{H}\mathord{\upharpoonright}A$ diagonalises $(A_{n})_{n<\omega}$ below $A$ if $B/n\leq A_{d_{A}(r_{n}(B))}$ for all $n<\omega$.
@@
    We say that $\mathcal{H}$ is selective if for all $A\in\mathcal{H}$ and $\leq$-decreasing chain $(A_{n})_{n<\omega}$ of elements in $\mathcal{H}$ below $A$, there exists a diagonalisation of $(A_{n})_{n<\omega}$ below $A$ in $\mathcal{H}$.
Smythe also proposed a notion of strong diagonalisation in Definition 4.1 of [23], followed by a notion of strong (p)-property. The following lemma shows that they are equivalent if $E$ is a vector space over a finite field. It is unclear if they are equivalent in case the underlying field is infinite.
@@
Suppose that $E$ is a vector space over a finite field. Let $\mathcal{H}\subseteq E^{[\infty]}$ be a semicoideal. The following are equivalent:
  1. (1)
    $\mathcal{H}$ is selective.
  2. (2)
    If $(A_{a})_{a\in E^{[<\infty]}}$ generates a filter in $\mathcal{H}$, then there exists some $B\in\mathcal{H}$ such that for all $n$, $B/n\leq A_{r_{n}(B)}$.
@@
[…]
$\mathcal{F}_{n}:=\{a\in E^{[<\infty]}\mathord{\upharpoonright}A:d_{A}(a)\leq n\}.$
Since the underlying field is finite, $\mathcal{F}_{n}$ is finite for all $n$. Thus, we may inductively define a $\leq$-decreasing sequence $(A_{n})_{n<\omega}$ in $\mathcal{H}$ such that $A_{n}\leq A_{a}$ for all $a\in\mathcal{F}_{n}$. Since $\mathcal{H}$ is selective, we may let $B\in\mathcal{H}$ be such that $B/n\leq A_{d_{A}(r_{n}(B))}\leq A_{r_{n}(B)}$ for all $n$. Then $B$ is the desired element in $\mathcal{H}$. ∎
[…]
We conclude the section by examining the relationship among the selectivity property, the (p)-property and the (q)-property.
Lemma 3.10.
Let $\mathcal{H}\subseteq E^{[\infty]}$ be a (q)-semicoideal. Let $(A_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}$. If $B\in\mathcal{H}\mathord{\upharpoonright}A_{0}$ weakly diagonalises $(A_{n})_{n<\omega}$, then there exists some $C\in\mathcal{H}\mathord{\upharpoonright}B$ which diagonalises $(A_{n})_{n<\omega}$ below $A_{0}$.
Proof.
Let $A:=A_{0}$, and let $B\in\mathcal{H}\mathord{\upharpoonright}A$ weakly diagonalise $(A_{n})_{n<\omega}$. Let $(I_{m})_{m<\omega}$ be any increasing sequence of intervals such that $B/\max(I_{m})\leq A_{d_{A}(r_{\min(I_{m})-1}(B))}$ for all $m$. By the (q)-property, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}B$ such that for all $n$, there exists some $m$ such that $\langle r_{n}(C)\rangle\subseteq\langle r_{\min(I_{m})-1}(B)\rangle$ and $C/n\leq B/\max(I_{m})$. The first property implies that $d_{A}(r_{n}(C))\leq d_{A}(r_{\min(I_{m})}(B))$, so by the second property:
@@
so $C$ diagonalises $(A_{n})_{n<\omega}$ below $A_{0}$. ∎
@@
$\langle r_{n}(B)\rangle\subseteq\langle r_{k}(A)\rangle\subseteq\langle r_{\min(I_{m(k)-1})}(A)\rangle,$
@@
4. Almost disjoint families of subspaces
@@ 4.1. Uncountability of almost disjoint families
We begin with the most fundamental question of the theory of mad families — the uncountability of infinite mad families. Consider the following two cardinals:
@@
It is clear that $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}\leq\mathfrak{c}$. By Proposition 2.5 of [24], $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ (in fact, $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$), so $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ is a cardinal invariant. One may ask if the same holds for $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ — that is, if $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ is true. Smythe’s proof of $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ heavily relies on the following lemma:
Lemma 4.3 (Lemma 2.3, [24]).
Let $A\in E^{[\infty]}$ be an infinite block sequence, and let $a\in E^{[<\infty]}$ be a finite block sequence. Then there exists some $N<\omega$, that depends only on $A$ and $\max(\operatorname{supp}(a))$, such that for all $x>N$ for which $x\notin\langle A\rangle$,
@@
He remarked that removing the “block” requirement from this lemma seems difficult. $\aleph_{1}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ was also stated as Proposition 3.1 of [12], but Kolman further assumed that in the countable almost disjoint family of subspaces $\{V_{n}:n<\omega\}$, $V_{n}\cap\sum_{i\in I}V_{i}$ is finite-dimensional for all finite $n\notin I\subseteq\omega$. This is generally not satisfied by families of subspaces that are pairwise almost disjoint. The author raised this question on MathOverflow, of which user527492 provided an affirmative answer (see [29]).
@@
Let $U,V,W\subseteq E$ be subspaces such that $V$ and $W$ are almost disjoint, and $U$ is finite-dimensional. Then $U+V$ and $W$ are almost disjoint.
@@
Define a linear map $\phi:V\times W\to E$ by stipulating that $\phi(v,w):=w-v$. Let $\rho:E\to E/U$ be the canonical quotient map, and let $\psi:=\rho\circ\phi$. Note that:
@@
Let $\pi:V\times W\to W$ be the projection map. Observe that for all $w\in W$, we have that:
@@
We shall define the subspace $W=\langle x_{n}:n<\omega\rangle$ by defining $x_{n}$ inductively as follows: Let $x_{0}\in V_{0}$ be any non-zero vector. Now assume that we have defined a linearly independent set $\{x_{i}:i<n\}$ such that for all $k<n$, we have that:
$\langle x_{i}:i<n\rangle\cap V_{k}\subseteq\langle x_{i}:i\leq k\rangle.$
We wish to find some vector $x_{n}\in V_{n}$, linearly independent from $\{x_{i}:i<n\}$, such that for all $k<n$, we have that:
$\langle x_{i}:i\leq n\rangle\cap V_{k}\subseteq\langle x_{i}:i\leq k\rangle.$
If we can complete this induction step, then we are done as we have that $W\cap V_{n}\subseteq\langle x_{0},\dots,x_{n}\rangle$ is finite-dimensional for all $n$.
For each $k<n$, let $U_{k}:=V_{k}+\langle x_{i}:i<n\rangle$. By Claim 5, $V_{n}\cap U_{k}$ is of finite dimension. Since no infinite-dimensional vector space is a finite union of finite-dimensional subspaces, we have that:
@@
so we may let $x_{n}$ be any vector in this set. We shall show that this vector $x_{n}$ works. We first note that $\{x_{i}:i\leq n\}$ is linearly independent, as $\langle x_{i}:i<n\rangle\subseteq U_{0}$ and $x_{n}\notin U_{0}$. Now let $k<n$ and $x\in\langle x_{i}:i\leq n\rangle\cap V_{k}$, so we have that $x=\sum_{i\leq n}\xi_{i}x_{i}\in V_{k}$. If $\xi_{n}\neq 0$, then we have that:
$x_{n}=\xi_{n}^{-1}\left(x-\sum_{i<n}\xi_{i}x_{i}\right)\in V_{n}\cap U_{k},$
contradicting the choice of $x_{n}$. Therefore, $\xi_{n}=0$, which implies that:
$x\in\langle x_{i}:i<n\rangle\cap V_{k}=\langle x_{i}:i\leq k\rangle\cap V_{k},$
@@
Smythe asked in [24] whether $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}=\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ holds in $\mathsf{ZFC}$, and this problem remains open to the author’s knowledge. In the same paper, Smythe showed that every finite almost disjoint family of block subspaces of cardinality greater than $1$ is not maximal. We extend this fact to arbitrary subspaces.
@@
Let $\mathcal{A}$ be a finite almost disjoint family of subspaces of $E$ such that $|\mathcal{A}|>1$. Then $\mathcal{A}$ is not maximal.
@@
Let $\mathcal{A}=\{V_{k}:k<n\}$ be a finite almost disjoint family of subspaces of $E$.
@@
We first assume that for all $j\neq k$, $V_{j}\cap V_{k}=\{0\}$. Let $\mathbb{F}$ be the field that $E$ is defined over. If $\mathbb{F}$ is infinite, then we are done, as any subspace over an infinite field is not a finite union of proper subspaces. Assume that $|\mathbb{F}|=N<\infty$, and suppose for a contradiction that $E=\bigcup_{k<n}V_{k}$. Let $M$ be some large number (which we shall fix later), and let $W:=\langle e_{n}:n<M\rangle$. Let $U_{k}:=W\cap V_{k}$, and by taking $M$ large enough we may assume that $U_{k}$ is non-trivial for all $k$. Note that $W=\bigcup_{k<n}U_{k}$.